OpenStudy (anonymous):

Which ratio forms a proportion with 12/10 3/5 60/50 26/20 5/6

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hello, my name is \(\Large\bf\color{#872af0}{Ⓐ}~\color{#d8b700}{Compassionate}\) In order to figure this out, you want to take each term in your options, and cross multiply it. If the numerators are equal, then that is a solution!

OpenStudy (anonymous):

so A?

OpenStudy (compassionate):

\[\frac{ 12 }{ 10 } *\frac{ 3 }{ 5 } = 60 -- 30. \] The numerators are 60 and 30. C'mon, you're not even tryin'! This is really easy.

OpenStudy (anonymous):

but i did it and i got 60 and 60

OpenStudy (compassionate):

If you cross multiply, 12 * 5 = 60 And 3 * 10 = 30 It's not A.

OpenStudy (anonymous):

ohhh B

OpenStudy (compassionate):

Correct! 60 and 600! Good job! OpenStudy On ~~